OpenStudy (anonymous):

Find Sn for the given geometric series a1=0.25, a5=446.265625, r=6.5

OpenStudy (anonymous):

common ratio r = 4th root (446.265625 / 0.25) = 6.5 Sn = 0.25 * 6.5^n - 1 -------- 6.5 - 1

OpenStudy (anonymous):

r = 4th root as a5 = a1r^4

OpenStudy (anonymous):

and, of course , a1 r^4 / a1 = r^4

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

thanks for the help joey

OpenStudy (anonymous):

yw